OpenStudy (anonymous):

simplify 5^3y^2/5^2y^-2 a. 5y^7 b. 5/y^3 c. 5y^3 d. 5^5y^7 HELP

jhonyy9 (jhonyy9):

do you know the property of exponents ? like x^(-1) = 1/x or x^a /x^b = x^(a-b) ?

jhonyy9 (jhonyy9):

so using these above wrote how many will be 5^3 --- = ? 5^2

jhonyy9 (jhonyy9):

us the last formula what i have wrote above